Operation Summary for November 24, 2010 - November 29, 2010 |
The Integrated Luminosity for the period from 11/22/10 to 11/29/10 was 57.5 Pb-1 (inverse picobarns). NuMI reported receiving 8.4E18 protons on target during this same period. Wednesday Nov 24 The day shift began with the Tevatron (TeV) in shot setup, with Operations monitoring the Recycler stash (127.4E10), with the Antiproton Source (Pbar) stacking (32.3E10), with Linac, Booster, and Main Injector (MI) operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) waiting for the next store. During the day shift, Operations established store 8296 at 8:32 AM with an average initial luminosity (AIL) of 303E30. Operations reported some Pbar downtime due to an access into the rings to repair a Programmable Logic Controller (PLC03). During the evening shift, Operations reported that the NuMI beam permit dropped several times due to a bad beam position (the vertical position at 607 was too high). Operators tuned to bring the position back into tolerance. Thursday Nov 25 The midnight shift began with Operations monitoring store 8296 (68.64E30) and the stash (342.4E10), with Pbar stacking (10E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Happy Thanksgiving Day! Operations terminated store 8296 at 1:18 AM, put the TeV into shot setup, and established store 8297 at 2:33 AM with an AIL of 186.8E30. Operations reported that TeV beam tune was off and landed too low, which caused the low initial luminosity. During the day shift, Operations terminated store 8297 at 10:36 AM. Operators investigated why the antiprotons moved out of their normal position for store 8297; this lead to some tuning and angle changes. Operations put the TeV into shot setup and established store 8300 at 2:55 PM with an AIL of 227.2E30. During the evening shift, operators resolved some FIRUS issues left over from the day shift. Operators tuned MI to improve beam positions with limited success. Operations reported that a Booster RF station (BRF12) failed; operators brought BRF7 online in its place. Friday Nov 26 The midnight shift began with Operations monitoring store 8300 (77.49E30) and the stash (247.3E10), with Pbar stacking (9.5E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations terminated store 8300 at 5:52 AM and put the TeV into shot setup, but the TeV quenched at TeV sector D1 and the D0 low beta magnets during the squeeze. Operator investigated and began the cryo system recovery. Operators reported that the regulation for a NuMI quadrupole power supply (E:QD113) failed and pulled the NuMI beam permit. Operations called in EE Support personnel who replaced a regulator card and Operations resumed sending beam to NuMI. During the day shift, operators had the TeV cryo system recovered by noon, but the quench investigation continued. Operator prepared D0 for a one-hour collision hall access. During the evening shift, Operations put the TeV into shot setup at 5:30 PM and established store 8307 at 6:35 PM with an AIL of 221E30. Saturday Nov 27 The midnight shift began with Operations monitoring store 8307 (95.63E30) and the stash (183.7E10), with Pbar stacking (25.8E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations terminated store 8307 at 3:20 AM, put the TeV into shot setup, and established store 8308 at 4:50 AM with an AIL of 269E30. During the day shift, Operations terminated store 8308 at 2 PM. Operators prepared D0 for a 1.5-hour access. During the evening shift, Operations put the TeV into shot setup at 5 PM and established store 8311 at 6:10 PM with an AIL of 293E30. Sunday Nov 28 The midnight shift began with Operations monitoring store 8311 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations reported that a TeV kicker (E48) tripped off due to low insulating gas (SF6); operators filled the system. The antiproton sequencer hung up once during transfers for loss of .4 mA. Operations reported that a Pbar Accumulator RF station (ARF1) tripped off twice. Operations terminated store 8311 at 5:50 AM and operators prepared D0 for a collision hall access. They came out of access at 6:40 AM. Operations began a low beta check and TeV orbit check. During the day shift, Operations put the TeV into shot setup at 8:15 AM and established store 8314 at 9:15 PM with an AIL of 308E30. During the evening shift, Operations terminated store 8314 at 11:39 PM and put the TeV into shot setup at 11:50 PM. Monday Nov 29 The midnight shift began with the TeV in shot setup, with Operations monitoring the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations established store 8315 at 1:05 AM with an AIL of 318E30. Operations reported that a Linac RF station (LRF2) tripped off numerous times. The Plans for Monday The plans for today are to stack, stash, and store. The Run Coordinator will allow MI personnel to conduct studies. |
|
More Information Comments and Suggestions |
| last modified 11/29/2010 email Fermilab |
| Security, Privacy, Legal | |